# Multiplexity amplifies geometry in networks

### v1

@@ Appendix B Geometric Multiplex Model
×$\times\left(\varphi_{1}^{1/(1-\nu)}+\varphi_{2}^{1/(1-\nu)}\right)^{-\nu}\frac{\varphi_{1}^{\nu/(1-\nu)}\kappa_{0}^{(1)}\left(\kappa_{i}^{(1)}\right)^{\gamma}_{1}}{\kappa_{0}^{(1)}\left(\kappa_{i}^{(1)}\right)^{\gamma}_{1}-\kappa_{0}^{(1)}\kappa_{i}^{(1)}},$ (5)
@@
If more layers are present this process is continued, generating the hidden variables in layer $l$ on the basis of those in layer $l-1$. In principle one could choose to vary $g$ and $\nu$ parameters, leading to some pairs of layers that are more strongly correlated than others. In this work we choose them identical between all layers.

### v2

@@ Acknowledgements.
Data Availability
The sources of the real data analyzed in this paper can be found in App. A. The corresponding results in tabular form are provided in the Supplementary Information [30]. The implementation and outputs of the numerical simulations performed in this study are available from the authors upon reasonable request.
@@ Appendix A Details real networks
  • Malaria [37]: The multiplex of recombinant antigen genes from the human Malaria parasite Plasmodium falciparum, where nodes are var genes encoding for an antigen protein expressed on the surface of the infected red blood cell. These var genes are capable of interchanging bits of genetic information, thus creating a vast amount of slightly different genes that in turn create highly varying antigens that are hard for the human immune system to detect. Two nodes are connected if they share a substring of nucleotides of significant length, indicating that an interchange of genetic material between the two nodes has taken place. The different layers represent different highly variable regions (HRVs).
@@ Appendix B Geometric Multiplex Model
$\times\left(\varphi_{1}^{1/(1-\nu)}+\varphi_{2}^{1/(1-\nu)}\right)^{-\nu}\frac{\varphi_{1}^{\nu/(1-\nu)}\kappa_{0}^{(1)}\left(\kappa_{i}^{(1)}\right)^{\gamma}_{1}}{\kappa_{0}^{(1)}\left(\kappa_{i}^{(1)}\right)^{\gamma}_{1}-\kappa_{0}^{(1)}\kappa_{i}^{(1)}},$ (5)
@@
If more layers are present this process is continued, generating the hidden variables in layer $l$ on the basis of those in layer l1l-1. In principle one could choose to vary $g$ and $\nu$ parameters, leading to some pairs of layers that are more strongly correlated than others. In this work we choose them identical between all layers.